OpenStudy (anonymous):

if x is a positive acute angle and cos x= 3/5, find the value of sin x

OpenStudy (ranga):

Use the same method as before.

OpenStudy (ranga):

cos(x) = adjacent / hypotenuse. So I drew a right triangle, marked one of the angles as x, used the definition of cos(x) as adjacent / hypotenuse and made the adjacent 3 and the hypotenuse 5 because cos(x) = 3/5. Use Pythagoras theorem to find AB. sin(x) - opposite / hypotenuse = AB / AC = ? / 5

OpenStudy (anonymous):

5^2+3^2=25+9=34 then what I do

OpenStudy (ranga):

In Pythagoras theorem, hypotenuse squared = sum of the squares of the other two sides. And hypotenuse is the side opposite the 90 degree angle. Here AC is the hypotenuse. So AC^2 = AB^2 + BC^2 AB^2 = AC^2 - BC^2 = 5^2 - 3^2 = ? AB = ?

OpenStudy (anonymous):

5.83095189485

OpenStudy (ranga):

No. You are still going by the wrong equation you wrote. If you look at my previous reply I corrected the error. You have to subtract 3^2 from 5^2 not add.

OpenStudy (anonymous):

Oh sorry i did not see the minus. So that mean its 14

OpenStudy (ranga):

AB^2 = 5^2 - 3^2 = 25 - 9 = 16 AB = 4 sin(x) = opposite / hypotenuse = AB / AC = 4 / 5

OpenStudy (anonymous):

Thank. You

OpenStudy (ranga):

You are welcome.